# How Easily do Irrelevant Inputs Skew the Responses of Large Language Models?

### v1

@@ Abstract
By leveraging the retrieval of information from external knowledge databases, Large Language Models (LLMs) exhibit enhanced capabilities for accomplishing many knowledge-intensive tasks. However, due to the inherent flaws of current retrieval systems, there might exist irrelevant information within those retrieving top-ranked passages. In this work, we present a comprehensive investigation into the robustness of LLMs to different types of irrelevant information under various conditions. We initially introduce a framework to construct high-quality irrelevant information that ranges from semantically unrelated, partially related, and related to questions. Furthermore, our analysis demonstrates that the constructed irrelevant information not only scores highly on similarity metrics, being highly retrieved by existing systems, but also bears semantic connections to the context. Our investigation reveals that current LLMs still face challenges in discriminating highly semantically related information and can be easily distracted by these irrelevant yet misleading contents. Besides, we also find that current solutions for handling irrelevant information have limitations in improving the robustness of LLMs to such distractions. Resources are available at https://github.com/Di-viner/LLM-Robustness-to-Irrelevant-Information.
@@ 2.2 Robustness to Irrelevant Information
Robustness, which refers to a system’s stability when confronted with unexpected inputs (Chang et al., 2023), has been extensively evaluated in previous studies on LLMs(Zhu et al., 2023; Chen et al., 2024). Given its potential to significantly impact model performance, irrelevant information has also attracted attention in the community (Shi et al., 2023a). Prior studies (Shi et al., 2023a; Wu et al., 2024) add specific instruction into prompts, enabling LLMs to better solve math word problems by automatically verifying the irrelevant content within problem descriptions. This approach can be combined with Chain-of-Thought (CoT) prompting methods (Wei et al., 2022; Kojima et al., 2022). However, these investigations primarily focus on irrelevant problem descriptions in arithmetic reasoning. In contrast, the challenge of irrelevant information in RAG applications arises more often from retrieved passages. Previous studies often classify low-ranked passages, random passages, and top-ranked passages without ground truth answers as irrelevant information (Yoran et al., 2023; Wang et al., 2023; Yu et al., 2023; Chen et al., 2024). Nonetheless, current advanced RAG systems may effectively filter out such content (Askari et al., 2023). In the real-world scenario, however, semantically related yet irrelevant information, which is highly likely to be retrieved by current systems, remains a challenge. To bridge this gap, our work meticulously constructs high-quality irrelevant information and offers a comprehensive analysis of LLM performance across various scenarios. This method enhances our understanding of LLMs’ interactions with irrelevant information, thereby providing valuable insights for improving the efficiency and effectiveness of RAG systems.
@@ 3.2 Parametric Memory Elicitation
3.3 Graded Irrelevant Information
Previous research has shown that LLMs can be easily distracted by irrelevant information, where even information with no relation to the topics of the questions they address can mislead LLMs (Shi et al., 2023a). However, there is a lack of detailed analysis concerning the degree of semantic relevance of irrelevant information that affects the performance of LLMs. To address this gap, we introduce a framework for categorizing irrelevant information into three graded levels, aiming to explore its impact in depth. Specifically, as shown in Figure 2 we define three distinct levels of irrelevant information: Unrelated Information, Partially Related Information, and Related Information.
@@ 3.4 Irrelevant Information Construction
To create realistic irrelevant information that is likely to be retrieved in real-world applications, we utilize Wikipedia, one of the largest databases, as our source. Following  Mallen et al. (2023), we use a dense retriever, the Contriever model (Gautier et al., 2022), to retrieve the Top 10 passages corresponding to each query. These passages are then used as candidate segments for constructing irrelevant information. Please refer to Appendix A.2 for more processing details. In subsequent parts of this subsection, we use the question “In what city was Julius Erving born?” as an illustrative example to clarify our construction strategy. Accordingly, we present the corresponding (subj, relationship, obj) triples of this question as (Julius Erving, place of birth, New York City).
@@
A) Parametric Memory: The city of East Meadow is located in Nassau County on Long Island, New York. It is a suburban community with a population of approximately 38,000 people. Julius Erving, also known as D̈r. J,ẅas born there on February 22, 1950. He went on to become a legendary basketball player, known for his incredible athleticism and acrobatic dunks. Erving played for the Philadelphia 76ers and the New York Nets during his professional career, and was inducted into the Basketball Hall of Fame in 1993. B) Unrelated Information: God’s mind as shadowed in the workings of the minds of men. Young ladies, if this degree has such meaning for your brothers, what meaning has it for you. Bonaparte lived in a townhouse in the north Baltimore neighborhood of Mount Vernon-Belvedere and had a country estate in suburban Baltimore County, Maryland, which surrounds the city on the west, north and east. His home, Bella Vista, was designed by the architects James Bosley Noel Wyatt, and William G. Nolting, in the prominent… C) Partially Related Information: they married in 2008. Erving has fathered nine children in total. Julius Erving Julius Winfield Erving II (born February 22, 1950), commonly known by the nickname Dr. J, is an American retired basketball player who helped popularize a modern style of play that… Baltimore is the most populous city in the U.S. state of Maryland. With a population of 585,708 at the 2020 census, it is the 30th-most populous city in the United States. Baltimore was designated an independent city by…
@@
D) Related Information - Misleading Linkage: During his career, Julius Erving played a memorable game in Baltimore, where he dazzled the crowd with his exceptional skills. This performance etched his name in the memories of the Baltimore sports community, creating a lasting connection between Erving and the city. Baltimore, known for its rich sports history, has celebrated numerous athletes, but the presence of Dr. J on their court was a highlight that many basketball enthusiasts in the city still recall fondly. E) Related Information - Common Characteristics: Julius Erving, often known as Dr. J, shared a commonality with C. J. Bonaparte in their dedication to excellence within their respective fields. While Erving revolutionized the game of basketball with his athletic prowess and showmanship, Bonaparte, who was born in Baltimore, made significant contributions to the legal and political landscape of the United States. Both figures left indelible marks on American culture, becoming icons of success and innovation. F) Related Information - Fictional Anecdotes: There’s an interesting anecdote that ties Julius Erving to the legacy of C. J. Bonaparte, who was born in Baltimore. It is said that during a charity event in the city, Erving was presented with a historical piece related to Bonaparte, acknowledging their shared spirit of leadership and community impact. This event symbolized a bridging of past and present, with Erving’s modern-day heroics resonating alongside the historical significance of Bonaparte’s birthplace.
@@
Compared with “Partially Related Information”, “Related Information” is highly semantically related to the question yet does not aid in answering it. To develop “Related Information” of high quality, we utilize the triples formed during the “Partially Related Information” stage, introducing additional misleading connections between the subject (subj) and the incorrect object (obj’) or another subject (subj’). Specifically, we create three variants of “Related Information”: 1) Misleading Linkage:This variant focuses on reinforcing the connection between subj and obj’. In the example in Table 1, Julius Erving and Baltimore are connected through his presence on the court of the city, which enhances the potential for confusion. 2) Common Characteristics:This variant highlights similarities between subj and another subj’, where the latter is associated with obj’. In the example, common characteristics between Erving and Bonabarte are contributions within their respective fields, thus adding a misleading layer of similarity. 3) Fictional Anecdotes:This variant creates scenarios involving subj and subj’, incorporating creative but irrelevant details. In the example, Erving received a historical Bonaparte item in Baltimore, linking their leadership legacies. We utilize GPT-4 Turbo to generate natural language information based on the above-mentioned triples and misleading connections. Please refer to Appendix A.3 for more details concerning the scales of the datasets at each step.
@@ 3.6 Evaluation Metrics
  • Misrepresentation Ratio. This metric assesses the rate at which LLMs modify their responses due to the influence of irrelevant information, effectively measuring their propensity to be misled by irrelevant information.
@@
To facilitate the answer parsing, following Xie et al. (2023), we adopt the multiple-choice QA format as the primary experimental framework to streamline the process of answer evaluation. To be specific, we offer options including memory answers of LLMs, irrelevant answers generated by human-written templates based on the misleading object obj’, and uncertainty option (i.e., “I’m not sure.”). Options are shuffled before being presented to LLMs. We provide more details and examples in Appendix A.4. Furthermore, in section 4.3, we investigate the impact of question formats on the performance of LLMs.
@@ 4 Experiments and Analysis
4.1 Semantic Relevance
Highly semantically related information is more likely to distract LLMs. As shown in Table 2, with the exception of Llama2-7B, LLMs relatively rarely change their memory answers to irrelevant options when facing “Unrelated Information”, which is unrelated to the question. However, challenges arise with information that is semantically related to the question. Taking GPT-3.5 Turbo as an example, compared to “Unrelated Information”, there is a notable increase in the misrepresentation ratio for “Partially Related Information” and “Related Information”, as well as the uncertainty ratio decreases. This indicates that LLMs are more prone to being misled by information that is highly semantically related but irrelevant to the question. Other models also exhibit diminished performance under similar conditions, particularly Llama2-7B. This suggests that current LLMs still struggle with discriminating irrelevant yet highly semantically related information. From these observations, in order to develop a reliable RAG system, the issue of irrelevant information interference should be taken seriously.
@@ 4.4 Limitations of Current Solutions
CoT might cause over-reasoning due to misleading irrelevant information. Figure 5 demonstrates that employing the CoT approach negatively affects the performance of GPT-3.5 Turbo, particularly implying its tendency to over-reason when presented with misleading information. For a detailed example of how GPT-3.5 Turbo errors under the influence of irrelevant information, please refer to Appendix B.3. This exploration suggests that while CoT can enhance reasoning depth, its effectiveness varies by model and context, potentially amplifying the risk of distraction by irrelevant details in certain scenarios.
@@
Figure 5: Solution attempts in a more complicated scenario, with all information presented to GPT-3.5 Turbo. Specifically, “‘+CoT” indicates prompting the model to think step by step, ‘+instr.” adds an instruction “feel free to ignore irrelevant information” into the prompt, and “+ICL” refers to adding examples to guide LLMs in discerning irrelevant information.
@@ A.1 List of Relationship Types and Question Templates
We provide a full list of relationship types and question templates for each dataset in Table A.1.
@@ A.2 Details of Retrieval Results
A.3 Scale of Datasets
@@
A.4 Creating Options: Process and Illustration
@@ Appendix B Details of Experiments and Analysis
B.1 Manual Evaluation of GPT-3.5 Turbo Accuracy of Free-form Question Alignment
@@
B.2 Question Format Influence on Responses of LLMs
@@
B.3 Details of Limitation of Current Solution.
@@
In a standard scenario (i.e., Vanilla), GPT-3.5 Turbo relies on its parametric memory, identifying Åkerö Castle was founded by the Swedish nobleman and statesman, Axel Oxenstierna. In contrast, when prompted to think step by step, and based on its inference from the detail that Carl Gustaf was a prominent figure in Swedish noble history and had extensive involvement in the development of numerous estates and castles throughout Sweden, GPT-3.5 Turbo is misled into selecting the irrelevant answer (i.e., Åkerö Castle was founded by Carl Gustaf Wrangel). In addition, despite being instructed to ignore irrelevant information and provided with exemplars, GPT-3.5 Turbo fails to differentiate the irrelevant details, ultimately selecting the irrelevant answer.

### v2

@@ Abstract
By leveraging the retrieval of information from external knowledge databases, Large Language Models (LLMs) exhibit enhanced capabilities for accomplishing many knowledge-intensive tasks. However, due to the inherent flaws of current retrieval systems, there might exist irrelevant information within those retrieving top-ranked passages. In this work, we present a comprehensive investigation into the robustness of LLMs to different types of irrelevant information under various conditions. We initially introduce a framework to construct high-quality irrelevant information that ranges from semantically unrelated, partially related, and related to questions. Furthermore, our analysis demonstrates that the constructed irrelevant information not only scores highly on similarity metrics, being highly retrieved by existing systems, but also bears semantic connections to the context. Our investigation reveals that current LLMs still face challenges in discriminating highly semantically related information and can be easily distracted by these irrelevant yet misleading content. Besides, we also find that current solutions for handling irrelevant information have limitations in improving the robustness of LLMs to such distractions. All the resources are available on GitHub.
@@ 2.2 Robustness to Irrelevant Information
Robustness, which refers to a system’s stability when confronted with unexpected inputs (Chang et al., 2023), has been extensively evaluated in previous studies on LLMs (Zhu et al., 2023; Chen et al., 2024). Given its potential to significantly impact model performance, irrelevant information has also attracted attention in the community (Shi et al., 2023a). Prior studies (Shi et al., 2023a; Wu et al., 2024) add specific instruction into prompts, enabling LLMs to better solve math word problems by automatically verifying the irrelevant content within problem descriptions. This approach can be combined with Chain-of-Thought (CoT) prompting method (Wei et al., 2022; Kojima et al., 2022). However, these investigations primarily focus on irrelevant problem descriptions in arithmetic reasoning. In contrast, the challenge of irrelevant information in RAG applications arises more often from retrieved passages. Previous studies often classify low-ranked passages, random passages, and top-ranked passages without ground truth answers as irrelevant information (Yoran et al., 2023; Wang et al., 2023; Yu et al., 2023; Chen et al., 2024). Nonetheless, current advanced RAG systems may effectively filter out such content (Askari et al., 2023). In the real-world scenario, however, semantically related yet irrelevant information, which is highly likely to be retrieved by current systems, remains a challenge. To bridge this gap, our work meticulously constructs high-quality irrelevant information and offers a comprehensive analysis of LLMs’ performance across various scenarios. This method enhances our understanding of LLMs’ interactions with irrelevant information, thereby providing valuable insights for improving the efficiency and effectiveness of RAG systems.
@@ 3.2 Parametric Memory Elicitation
3.3 Information Categorization
In this work, as shown in Figure 2, we categorize information into “Relevant” and “Irrelevant”. Specifically, following previous research (Cuconasu et al., 2024), we treat information that leads to or entails an exact answer as relevant, regardless of its correctness. Thus, gold passage and parametric memory passing a series of checks are considered relevant. Unlike relevant information, which suggests an exact answer, irrelevant information cannot support a possible answer and contains entities that distract LLMs. We aim to investigate whether LLMs are robust to these distractions. Previous research has shown that LLMs can be easily distracted by irrelevant information, where even information with no relation to the topics of the questions can mislead them (Shi et al., 2023a). However, there is a lack of detailed analysis concerning the degree of semantic relevance of irrelevant information that affects the performance of LLMs. To address this gap, we introduce a framework for categorizing irrelevant information into three graded levels, aiming to explore its impact in depth. Specifically, we define three distinct levels of irrelevant information: Unrelated Information, Partially Related Information, and Related Information.
@@ 3.4 Irrelevant Information Construction
To construct realistic irrelevant information that is likely to be retrieved in real-world applications, we utilize Wikipedia, one of the largest databases, as our source. Following Mallen et al. (2023), we use a dense retriever, the Contriever model (Gautier et al., 2022), to retrieve the Top 10 passages corresponding to each query. These passages are then used as candidate segments for constructing irrelevant information. Please refer to Appendix A.2 for more processing details. In subsequent parts of this subsection, we use the question “In what city was Julius Erving born?” as an illustrative example to clarify our construction strategy. Accordingly, we present the corresponding (subj, relationship, obj) triples of this question as (Julius Erving, place of birth, New York City).
@@
A) Parametric Memory: The city of East Meadow is located in Nassau County on Long Island, New York. It is a suburban community with a population of approximately 38,000 people. Julius Erving, also known as D̈r. J,ẅas born there on February 22, 1950. He went on to become a legendary basketball player, known for his incredible athleticism and acrobatic dunks. Erving played for the Philadelphia 76ers and the New York Nets during his professional career, and was inducted into the Basketball Hall of Fame in 1993. B) Unrelated Information: God’s mind as shadowed in the workings of the minds of men. Young ladies, if this degree has such meaning for your brothers, what meaning has it for you. Bonaparte lived in a townhouse in the north Baltimore neighborhood of Mount Vernon-Belvedere and had a country estate in suburban Baltimore County, Maryland, which surrounds the city on the west, north and east. His home, Bella Vista, was designed by the architects James Bosley Noel Wyatt, and William G. Nolting, in the prominent… C) Partially Related Information: they married in 2008. Erving has fathered nine children in total. Julius Erving Julius Winfield Erving II (born February 22, 1950), commonly known by the nickname Dr. J, is an American retired basketball player who helped popularize a modern style of play that emphasizes leaping… Baltimore is the most populous city in the U.S. state of Maryland. With a population of 585,708 at the 2020 census, it is the 30th-most populous city in the United States. Baltimore was designated an independent city by…
@@
D) Related Information - Misleading Linkage: During his illustrious career, Julius Erving played a memorable game in Baltimore, where he dazzled the crowd with his exceptional skills. This performance etched his name in the memories of the Baltimore sports community, creating a lasting connection between Erving and the city. Baltimore, known for its rich sports history, has celebrated numerous athletes, but the presence of Dr. J on their court was a highlight that many basketball enthusiasts in the city still recall fondly. E) Related Information - Common Characteristics: Julius Erving, often known as Dr. J, shared a commonality with C. J. Bonaparte in their dedication to excellence within their respective fields. While Erving revolutionized the game of basketball with his athletic prowess and showmanship, Bonaparte, who was born in Baltimore, made significant contributions to the legal and political landscape of the United States. Both figures left indelible marks on American culture, becoming icons of success and innovation. F) Related Information - Fictional Anecdotes: There’s an interesting anecdote that ties Julius Erving to the legacy of C. J. Bonaparte, who was born in Baltimore. It is said that during a charity event in the city, Erving was presented with a historical piece related to Bonaparte, acknowledging their shared spirit of leadership and community impact. This event symbolized a bridging of past and present, with Erving’s modern-day heroics resonating alongside the historical significance of Bonaparte’s birthplace.
@@
Compared with “Partially Related Information”, “Related Information” is highly semantically related to the question yet does not aid in answering it. To develop “Related Information” of high quality, we utilize the triples formed during the “Partially Related Information” stage, introducing additional misleading connections between the subject (subj) and the incorrect object (obj’) or another subject (subj’). Specifically, we create three variants of “Related Information”: 1) Misleading Linkage:This variant focuses on reinforcing the connection between subj and obj’. In the example in Table 1, Julius Erving and Baltimore are connected through his presence on the court of the city, which enhances the potential for confusion. 2) Common Characteristics:This variant highlights similarities between subj and another subj’, where the latter is associated with obj’. In the example, common characteristics between Erving and Bonabarte are contributions within their respective fields, thus adding a misleading layer of similarity. 3) Fictional Anecdotes:This variant creates scenarios involving subj and subj’, incorporating creative but irrelevant details. In the example, Erving received a historical Bonaparte item in Baltimore, linking their leadership legacies. We utilize GPT-4 Turbo to generate natural language information based on the above-mentioned triples and misleading connections. Please refer to Appendix A.3 for more details on generating Related Information and Appendix A.4 for more details on the scales of the datasets at each step.
@@ 3.6 Evaluation Metrics
  • Misrepresentation Ratio. This metric assesses the rate at which LLMs modify their responses due to the influence of irrelevant information, effectively measuring their propensity to be misled by irrelevant information.  [^1] Whether LLMs rely on parametric memory or gold passages is not considered misrepresentation, as both types of information are deemed “relevant”.
@@
To facilitate the answer parsing, following Xie et al. (2023), we adopt the multiple-choice QA format as the primary experimental framework to streamline the process of answer evaluation. To be specific, we offer options including memory answers of LLMs, irrelevant answers generated by human-written templates based on the misleading object obj’, and uncertainty options (i.e., “I’m not sure.”). Options are shuffled before being presented to LLMs. We provide more details and examples in Appendix A.5. Furthermore, in section 4.3, we investigate the impact of question formats on the performance of LLMs.
@@ 4 Experiments and Analysis
4.1 Semantic Relevance
Highly semantically related information is more likely to distract LLMs. As shown in Table 2, with the exception of Llama2-7B, LLMs relatively rarely change their memory answers to irrelevant options when facing “Unrelated Information”, which is unrelated to the question. However, challenges arise with information that is semantically related to the question. Taking GPT-3.5 Turbo as an example, compared to “Unrelated Information”, there is a notable increase in the misrepresentation ratio for “Partially Related Information” and “Related Information”, as well as the uncertainty ratio decreases. This indicates that LLMs are more prone to being misled by information that is highly semantically related but irrelevant to the question. Other models also exhibit diminished performance under similar conditions, particularly Llama2-7B. This suggests that current LLMs still struggle with discriminating irrelevant yet highly semantically related information. Additionally, the discussion regarding the uncertainty ratio is detailed in Appendix B.1. From these observations, in order to develop a reliable RAG system, the issue of irrelevant information interference should be taken seriously.
@@ 4.3 Question Format
Figure 5: Solution attempts in a more complicated scenario, with all information presented to GPT-3.5 Turbo. Specifically, “+CoT” indicates prompting the model to think step by step, “+instr.” adds an instruction “feel free to ignore irrelevant information” into the prompt, and “+ICL” refers to adding examples to guide LLMs in discerning irrelevant information. For a discussion about the model-based fine-tuning method, please refer to Appendix B.5.
@@ 4.4 Limitations of Current Solutions
CoT might cause over-reasoning due to misleading irrelevant information. Figure 5 demonstrates that employing the CoT method negatively affects the performance of GPT-3.5 Turbo, particularly implying its tendency to over-reason when presented with misleading information and infer wrong answers due to incorrect reasoning chains. This observation aligns with previous work on math problems, known as over-reasoning (Hou et al., 2024; Chiang & Lee, 2024). For a detailed example of how GPT-3.5 Turbo errs under the influence of irrelevant information, please refer to Appendix B.4. This exploration suggests that while CoT can enhance reasoning depth, its effectiveness varies by model and context, potentially amplifying the risk of distraction by irrelevant details in certain scenarios.
@@ A.1 List of Relationship Types and Question Templates
We provide a full list of relationship types and question templates in Table A.1.
@@ A.2 Details of Retrieval Results
A.3 Related Information Generation
We provide details of the prompts for generating Related Information in Table A.2. To prevent the generation of false context, we add an instruction such as: “However, you are not allowed to claim or hint that ’Julius Erving was/was not born in Baltimore’.” In addition, we instruct GPT-4 Turbo to return a “null” response when it encounters difficulties.
You are tasked with creating a Question and Answer scenario. Your goal is to compose three distinct paragraphs, each designed to subtly mislead or distract from the actual answer to the question, ’In what city was Julius Erving born?’ These paragraphs should be crafted in a way that strongly connects Julius Erving with Baltimore. However, you are not allowed to claim or hint that ’Julius Erving was/was not born in Baltimore’. Your response is required in JSON format: { xx”evidence”: { xxxx”paragraph1”: ””, xxxx”paragraph2”: ””, xxxx…, xx} } The content in the ”paragraph” should be ”null” if it is hard to compose such a specific paragraph. 1. Find some common characteristics or a connection between Julius Erving and Charles Joseph Bonaparte, mentioning that Charles Joseph Bonaparte was born in Baltimore. 2. Identify a connection between Julius Erving and Baltimore. 3. Create an anecdote involving Julius Erving and Charles Joseph Bonaparte, ensuring that the information ”Charles Joseph Bonaparte was born in Baltimore.” is mentioned.
Table A.2: Prompts for generating Related Information.
A.4 Scale of Datasets
@@
A.5 Creating Options: Process and Illustration
@@ Appendix B Details of Experiments and Analysis
B.1 Discussion on Uncertainty Ratio
Table 2 demonstrates that as the semantic relevance of the information increases, the uncertainty ratio decreases. To explore this phenomenon in detail, we sample 100 data points from EntityQuestions to observe GPT-3.5 Turbo’s responses. The results are shown in Table B.5. We find that LLMs might choose the uncertainty option for security reasons, especially when facing Unrelated Information.
Specifically, in uncertain situations when facing Unrelated Information, 57 out of 73 responses explain “The given information doesn’t provide any details about the question”, while the rest simply state the option. This indicates that GPT-3.5 Turbo identifies Unrelated Information but expresses uncertainty for conservative reasons.
Unrelated PartRel. Related
Uncertainty Count 73 39 22
Table B.5: GPT-3.5 Turbo chooses uncertainty options when facing different information.
B.2 Manual Evaluation of GPT-3.5 Turbo Accuracy of Free-form Question Alignment
@@ B.1 Discussion on Uncertainty Ratio
B.3 Question Format Influence on Responses of LLMs
@@
B.4 Details of Limitation of Current Solution.
@@
In a standard scenario (i.e., Vanilla), GPT-3.5 Turbo relies on its parametric memory, identifying Åkerö Castle was founded by the Swedish nobleman and statesman, Axel Oxenstierna. In contrast, when prompted to think step by step, and based on its inference from the detail that Carl Gustaf Wrangel was a prominent figure in Swedish noble history and had extensive involvement in the development of numerous estates and castles throughout Sweden, GPT-3.5 Turbo is misled into selecting the irrelevant answer (i.e., Åkerö Castle was founded by Carl Gustaf Wrangel). In addition, despite being instructed to ignore irrelevant information and provided with exemplars, GPT-3.5 Turbo fails to differentiate the irrelevant details, ultimately selecting the irrelevant answer.
B.5 Results of Fine-tuning Method
In our paper, we focus on prompted-based methods to help discern irrelevant information. As fine-tuning is a proven method to enhance the robustness of LLMs to irrelevant information (Yoran et al., 2023), we additionally sample 100 data points to test how llama2-13b and llama2-13b-peft-nq-retrobust (+ft.) perform with constructed related information. As shown in Table B.6, the misrepresentation ratio (MR) indicates that the fine-tuned model shows improved robustness, but remains high, suggesting challenges still exist.
MR
llama2-13b 61
llama2-13b + ft. 25
Table B.6: Performance of Llama2-13b and fine-tuned models on 100 sampled data points.